Annual income: The mean annual income for people in a certain city in thousands of dollars) is 41, with a standard deviationPart 4 of 5 (d) Would it be unusual for the sample mean to be less than 33? Round the answer to at least four decimal places.
Answer #1

\mu = 41

\sigma = 35

n = 91

For sampling distribution of mean, P(\bar{x} < A) = P(Z < (A - \mu_\bar{x} )/\sigma_\bar{x})

\mu_\bar{x} = \mu = 41

\sigma_\bar{x} = ulo

= 35/V91

= 3.669

1) P(\bar{x} < 37) = P(Z < (37 - 41)/3.669)

= P(Z < -1.09)

= 0.1379

2) P(40 < \bar{x} < 45) = P(\bar{x} < 45) - P(\bar{x} < 40)

= P(Z < (45 - 41)/3.669) - P(Z < (40 - 41)/3.669)

= P(Z < 1.09) - P(Z < -0.27)

= 0.8621 - 0.3936

= 0.4685

3) Let the 30th percentile be T

P(\bar{x} < T) = 0.30

P(Z < (T - 41)/3.669) = 0.30

Take the value of Z corresponding to 0.30 from standard normal distribution table.

(T - 41)/3.669 = -0.52

T = 39.09

4) A value is unusual if the probability of occurrence is less than 0.05

P(\bar{x} < 33) = P(Z < (33 - 41)/3.669)

= P(Z < -2.18)

= 0.0146

It is unusual because the probability of the sample mean being less than 33 is 0.0146

5) P(X < 33) = P(Z < (33 - \mu )/\sigma)

= P(Z < (33 - 41)/35)

= P(Z < -0.23)

= 0.4090

It is not unusual because the probability of the sample mean being less than 33 is 0.4090